Use a calculator to graph the function. Then, using the graph, give three points on the graph of the inverse with the y-coordinates given.
f(x) = x^4 + 6x + 4, −1 ≤ x ≤ 1, y = −1,1
Y= -1 (X, Y)=
Y= 1 (X, Y)=

* I understand that the inverses are the original points reversed, so to solve the problem I would simply need to plug in each point given into the equation, above, arrive at the answer, and reverse the points in order to transform it into an inverse function. However, I’m not sure how −1 ≤ x ≤ 1 factors into the solution.

Answer by stanbon(75887)

Use a calculator to graph the function. Then, using the graph, give three points on the graph of the inverse with the y-coordinates given.
f(x) = x^4 + 6x + 4, −1 ≤ x ≤ 1, y = −1,1
If Y= -1 (X, Y) =
Solve -1 = x^4+6x+4 where -1<=x<=1
x^4+6x+5 = 0
(x^2+1)(x^2+5) = 0
x = 1 or x = -1
That gives you the points (1,-1) and (-1,-1)
The corresponding points of the inverse would be (-1,1) and (-1,-1)
--------
Y= 1 (X, Y)=
Solve 1=x^4+6x+4 where -1<=x<=1
x^4+6x+3 = 0
Pick a point on the graph where -1<=x<=1::
Example:: (0,3)
The corresponding point on the the inverse would be (3,0)
------------
Cheers,
Stan H.
-----------
* I understand that the inverses are the original points reversed, so to solve the problem I would simply need to plug in each point given into the equation, above, arrive at the answer, and reverse the points in order to transform it into an inverse function. However, I’m not sure how −1 ≤ x ≤ 1 factors into the solution.